# Beyond Trial-and-Error: Predicting User Abandonment After a Moderation Intervention

### v2

@@ VIII-C Leave-one-out cross-validation
Figure 5: Top-10 most important features for the hard and soft abandonment tasks. The ranking of the most important features is similar between the two tasks, with features such as trend_ext, part_ratio, and avg_id_attack dominating the ranking. All classes of features are represented in the top-10, with the exception of writing style features that do not seem to provide individually strong contributions to models predictions.
VIII-D Feature importance
@@
The $\text{score}_{L}(i,j)$ expresses the local contribution, or importance, of the feature $f_{i}$ for the model $M_{j}$. Next, we obtain a global score for the feature $f_{i}$ as the weighted mean of its local importance across all trained models:
@@
We use the F1 score on the positive class as the weighting factor $w_{j}$ of each model $M_{j}$, so that models that are better at detecting abandoning users are weighted more in Eq. (3). The $\text{score}_{G}(i)$ expresses the weighted contribution of the feature $f_{i}$ throughout all models that we trained and tested —hence the global label— thus representing a robust indicator of the actual importance of that feature for the task. As a final presentation step, we rank all features and assess their relative importance by normalizing their global scores so that the best performing feature has a normalized score of 1 and the others are rescaled proportionally.
@@
Figures 5(a) and 5(b) show the top-10 most important features according to their normalized global scores, respectively for the hard and soft abandonment tasks. The set of the most important features for both tasks is very similar, although the ranking is different and the features exhibit different relative importance. In both tasks, among the most informative features are the trend of comments in non-banned subreddits (trend_ext), the ratio between the number of non-banned and banned subreddits to which a user participated (part_ratio), and the average score of “identity attacks” in the user’s comments (avg_id_attack). For the hard abandonment task, in particular, the contribution of the trend_ext feature is almost double that of the second-best one. To this end we recall that the DT Trend baseline defined in Section VII-A is based on this exact feature, which explains the good performance of that baseline in the results of Table II. By observing the relative importance between the top and bottom features in Figures 5(a) and 5(b), we also derive further insights into the results of Table II. Specifically, we note that in both tasks the first few features provide a much larger relative contribution than the last ones in the top-10. This means that, apart from the first few features, all others, including those highly ranked according to our feature importance analysis, provide relatively few information to the models. This result resonates with that presented in Table II about the small number of features used by the majority of models. Finally, we note that three out of four classes of features are represented among the top-10 features shown in Figure 5. The only class of features that did not make it to the top-10 is the one conveying information the writing style of the users, which does not seem to provide much relevant information.
(a) Hard abandonment.
(b) Soft abandonment.
@@
Figure 6: Feature importance of each class of features, obtained by summing the contributions of the individual features in each class, normalized by the number of features in the class. The contribution of the different classes of features is largely the same in both tasks, with activity features providing the largest contribution, followed by relational and toxicity features. Writing style features provide overall marginal contributions.
To better assess the contribution of the different classes of features, rather than that of the individual features, we aggregate the contributions of all features based on their class. In detail, we sum the global importance scores of all features separately for each class, and we divide it by the number of features in the class, to obtain a class feature importance score. Lastly, we normalize the class scores so that the best class has a normalized score of 1 and the others are rescaled proportionally.
@@
Figure 6 shows the relative importance of the different classes of features for the two tasks. The ranking is the same in both tasks, with activity features providing the largest contribution. Relational features are the second-most informative class, providing about 80% of the contribution of activity features. Toxicity features obtain a similar score. Instead, writing style features are ranked last in both tasks and provide small contributions with respect to the other classes. This result confirms that shown in Figure 5, where no writing style feature was ranked among the top contributing features in either task. Overall, the results obtained for activity, relational, and toxicity features highlight the importance of user engagement, social dynamics, and toxic speech in predicting user abandonment after the investigated moderation intervention. Conversely, information about the writing style of the users does not appear to be as relevant.
IX Discussion
IX-A Detecting abandoning users
@@
In this study, we defined and tackled the novel task of predicting user abandonment following a massive deplatforming moderation intervention on Reddit [11]. In consideration of the novelty of the task —tackled here for the first time— our results are promising. For example, we achieved micro F1 $=0.800$ (macro F1 $=0.676$) when detecting users who immediately halted activity after the moderation intervention (hard abandonment), and micro F1 $=0.737$ (macro F1 $=0.660$) when detecting users who initially maintained some activity but who eventually left Reddit (soft abandonment). Albeit preliminary, these results demonstrate that the task of preemptively estimating the effects of future moderation interventions is indeed feasible. Adding to the encouraging findings are the results of our leave-one-out cross-validation (LOOCV) analysis, which underscore the absence of substantial performance degradation when applying our model to unseen data and shows that the model learned generalizable behavioral patterns. At the same time however, our study highlighted some of the challenges of this new task. Among them are the difficulties at forecasting complex and heterogeneous user reactions [18], a dire class imbalance [33], a partial and unreliable view of platform moderation actions [3], and the growing difficulties at accessing platform data [57], which culminate in the lack of extensive labeled datasets. The inherent challenges of predicting the effects of moderation interventions —exemplified by our promising yet preliminary results— imply that much work remains to be done before moderators and platform administrators will be able to leverage powerful and dependable tools for accurately estimating the effects of their actions. To this end, this study paves the way for future endeavors in this emerging area of content moderation, which proposes a new pathway for empowering moderators to optimize interventions based on desired outcomes, such as reducing toxicity while minimizing user attrition and churn. This is currently an as-yet unexplored frontier of moderation, which will open up once it achieves high performance in tasks such as the one proposed and tackled in this work.
IX-B Characterizing moderated users
@@
Although the performances of the models developed in this study are promising, there is much room for improvement. Model performance strongly depends on the informativeness of the provided features. In this initial study, we computed and tested an extensive and diverse set of 142 features, many of which were borrowed from recent literature on related tasks [58, 28, 44, 41]. Our set of features included information about the activity of the users before the moderation intervention, their participation in multiple communities, their relationships with other users, their use of toxic or otherwise aggressive speech, and their writing style. While some of the extracted features turned out to be powerful predictors in our models, a large share of the features that we evaluated was not particularly informative, which hindered model performance. This suggests that our task, proposed here for the first time, is substantially different from others previously addressed in related literature and for which those features provided robust performance. This finding underscores the importance of tailoring feature selection to the specific characteristics of the task, rather than relying on already-proposed features. Future efforts towards the prediction of moderation intervention effects should thus consider implementing and experimenting with a broader and even more diversified set of features. To this end, promising directions for future work involve the development of socio-psychological features. In fact, predicting the effects of moderation interventions essentially revolves around predicting user behavior. Moreover, it is known that socio-psychological characteristics influence online behaviors, including toxic and aggressive ones [18]. A recent demonstration of the latter can be found in those works who measured a significant improvement in hate speech detection tasks, when incorporating socio-psychological features [59]. By the same token, those features could provide valuable information also in the related task of predicting behavioral changes following a moderation intervention, and particularly for those interventions targeting hateful or toxic users.
IX-C Predicting effects: classification, quantification, regression
@@
IX-D Limitations
@@
This study focuses on a set of Reddit users who experienced a specific form of moderation intervention (i.e., multiple community bans), which may limit the generalizability of our findings. Each platform operates within a unique ecosystem with its own user demographics, interaction norms, and moderation policies. Therefore, in spite of our efforts at estimating the generalizability of our results, findings derived from a single platform and for a single intervention may not be fully applicable to other platforms or interventions. Moreover, we focused on a subset of all moderated users: those who were particularly active in the banned communities. As such, our models might exhibit decreased performance if applied to users with markedly different characteristics. Additionally, the specific definitions of hard and soft abandonment that we proposed here might represent a further limitation, given that alternative definitions may yield different interpretations of user behavior and corresponding ground truths for the machine learning tasks. Another limitation of our work stems from the set of features with which we experimented. Despite implementing a relatively large number of features that convey diverse information, many important dimensions of user behavior remain unexplored for this task. Encoding those dimensions in effective machine learning features would boost the performance of the trained models, allowing better results. By the same token, in this first study on predicting the effect of a moderation intervention, we focused on detecting abandoning users. However, effects of a moderation intervention can be defined and investigated in multiple other ways that do not necessarily involve user abandonment, but rather the degree of toxic speech on the platform after the intervention, the extent of polarization, the reliance on factual news, and more. The predictability of these and other moderation effects remains to be assessed.
X Conclusions

### v3

@@ 8.3 Leave-one-out cross-validation
Figure 7: Top-10 most important features for the hard and soft abandonment tasks according to SHAP (top row) and PI (bottom row). The ranking of the most important features is similar between the two tasks, with features such as trend_ext, part_ratio, and avg_id_attack dominating the ranking. All classes of features are extensively represented in the top-10, with the exception of writing style features that do not seem to provide individually strong contributions to model predictions.
8.4 Feature importance
Our next analysis involves estimating the importance of the individual features as well as of the classes of features that we implemented, for both the hard and soft abandonment tasks. This analysis complements our previous results by providing additional information for interpreting model predictions and by highlighting which features are most predictive of user abandonment following a moderation intervention. Not only does such analysis enhance the interpretability of our models, but it also provides moderators with valuable insights into the key factors of user abandonment. By understanding which features contribute the most to the predicted outcomes, moderators would be able to identify patterns and refine their moderation strategies accordingly. In general, transparency can help moderators make better informed decisions, balancing rule enforcement with user retention, and ensuring that moderation remains effective and fair.
We rely on SHapley Additive exPlanations (SHAP) and Permutation Importance (PI) as the building blocks of our feature importance analysis. PI was introduced in Section 8.2. Instead, SHAP is a method based on cooperative game theory that is widely used for interpreting the output of machine learning models and for estimating the contribution of individual features to a model’s predictions [78, 60]. Leveraging two different post-hoc explainers allows us to gain a deeper understanding of the importance of the implemented features for this task.
@@ 8.4 Feature importance
The $\text{score}_{L}(i,j)$ expresses the local contribution —or importance— of the feature $f_{i}$ for the model $M_{j}$. Next, we obtain a global score for the feature $f_{i}$ as the weighted mean of its local importance across all trained models:
@@
We use the F1 score on the positive class as the weighting factor $w_{j}$ of each model $M_{j}$, so that models that are better at detecting abandoning users are weighted more in Eq. (3). The $\text{score}_{G}(i)$ expresses the weighted contribution of the feature $f_{i}$ throughout all models that we trained and tested —hence the global label— thus representing a robust indicator of the actual importance of that feature for the task. Contrarily to SHAP, PI estimates feature importance for whole models, rather than for individual data points. Thus, when computing feature importance with PI we directly apply Eq. (3) to obtain the global feature importance scores. Before that, we remove correlated features (i.e., those with Pearson $r>|0.7|$), as already done in Section 8.2, to avoid possible misleading results. As a final presentation step, we rank all features and assess their relative importance by normalizing their global scores so that the best performing feature has a normalized score of 1 and the others are rescaled proportionally.
@@
Figure 7 shows the top-10 most important features according to their normalized global scores, for the hard and soft abandonment tasks, with feature importance computed via SHAP and PI. The set of the most important features for both tasks and explainers is very similar, although the ranking is different and the features exhibit different relative importance. In both tasks, among the most informative features are the trend of comments in non-banned subreddits (trend_ext), the ratio between the number of non-banned and banned subreddits to which a user participated (part_ratio), the average score of “identity attacks” in the user’s comments (avg_id_attack), and the average time passed between the post of two subsequent comments (time_diff). For both tasks and explainers, the single most informative feature is by far trend_ext. To this end we recall that the DT Trend baseline defined in Section 7.1 is based on this exact feature, which explains the good performance of that baseline in Table 2. We also note that in both tasks the first few features provide a much larger relative contribution than the last ones in the top-10. This means that, apart from the first few features, all others, including those highly ranked according to our feature importance analysis, provide relatively few information to the models. This result resonates with that presented in Table 2 about the small number of features used by non-ensemble models. Finally, we note that three out of four classes of features are extensively represented among the top-10 features shown in Figure 7. The only class of features that did not make it to the top-10 —except for the 9th ranked features according to PI for the soft abandonment task— is the one conveying information the writing style of the users, which does not seem to provide much relevant information.
(a) Task: hard abandonment. Feature importance: SHAP.
(b) Task: soft abandonment. Feature importance: SHAP.
(c) Task: hard abandonment. Feature importance: PI.
(d) Task: soft abandonment. Feature importance: PI.
@@
Figure 8: Feature importance for the hard and soft abandonment tasks according to SHAP (top row) and PI (bottom row). The importance is calculated for each class of features by summing the contributions of individual features in each class, normalized by the number of features in that class. The contributions of the different classes of features are largely the same across the two tasks, with activity features providing the largest contribution, followed by relational and toxicity features. Writing style features provide overall marginal contributions.
@@
To better assess the contribution of the different classes of features, rather than that of the individual features, we aggregate the contributions of all features based on their class. In detail, we sum the global importance scores of all features separately for each class, and we divide it by the number of features in the class, to obtain a class feature importance score. Lastly, we normalize the class scores so that the best class has a normalized score of 1 and the others are rescaled proportionally. Figure 8 shows the relative importance of the different classes of features for the two tasks according to SHAP and PI. The ranking is the same for both tasks and explainers, with activity features providing the largest contribution. Relational features are the second-most informative class, providing about 80% of the contribution of activity features. Toxicity features obtain a similar score. Instead, writing style features are ranked last in both tasks and provide small contributions with respect to the other classes. This result confirms that shown in Figure 7, where no writing style feature was ranked among the top contributing features in either task. Overall, the results obtained for activity, relational, and toxicity features highlight the importance of user engagement, social dynamics, and toxic speech in predicting user abandonment after the investigated moderation intervention. Conversely, information about the writing style of the users does not appear to be as relevant.
8.5 Class imbalance
positive class overall
task rebalancing strategy p precision recall F1 AUC micro F1
hard abandonment 50 0.002 0.500 0.004 0.850 0.738
Resampling 50 0.520 0.373 0.434 0.781 0.797
Cost-Sensitive Learning 60 0.679 0.303 0.419 0.719 0.776
Resampling + Cost-Sensitive Learning 40 0.689 0.303 0.421 0.717 0.769
Ensemble Learning 60 0.107 0.542 0.179 0.853 0.770
Ensemble Learning + Resampling 70 0.440 0.324 0.373 0.779 0.742
Ensemble Learning + Cost-sensitive Learning 20 0.107 0.464 0.174 0.848 0.747
soft abandonment 40 0.024 0.583 0.046 0.732 0.779
Resampling 50 0.484 0.513 0.499 0.728 0.737
Cost-sensitive Learning 60 0.613 0.440 0.513 0.685 0.721
Resampling + Cost-Sensitive Learning 60 0.639 0.449 0.528 0.691 0.726
Ensemble Learning 80 0.161 0.592 0.253 0.744 0.719
Ensemble Learning + Resampling 80 0.485 0.486 0.485 0.723 0.725
Ensemble Learning + Cost-sensitive Learning 10 0.047 0.569 0.086 0.733 0.583
$\dagger$ Resampling = SMOTE + Random Undersampling
Table 6: Support Vector Machine (SVM) classification results in the hard abandonment (top rows) and soft abandonment (bottom rows) tasks, obtained by using different rebalancing strategies. Column p reports the number of features used by each model. For each task, the best result in each evaluation metric is shown in bold and the second-best is underlined.
Motivated by the observed degrading effect of class imbalance, we trained multiple Support Vector Machines (SVM) to experiment with a few different rebalancing strategies, used both in isolation and in combination. Among the strategies that we considered are SMOTE, Random Undersampling, Cost-Sensitive Learning, and Ensemble Learning. In addition, we also trained a classifier on the natural class distribution, without carrying out any rebalancing. We discussed the use of SMOTE and Random Undersampling in Section 7.2.4. Cost-Sensitive Learning balances the class weights by adjusting the loss function so as to increase the penalty for misclassifying the minority class, thus making the model pay more attention to correctly identifying those minority instances. Finally, the Ensemble Learning strategy implements a majority-voting ensemble by training ten classifiers on the dataset with a non-deterministic process, and then combining their outputs via majority voting.
Table 6 reports the results of these analyses. For both tasks, the results clearly show that the SVM without rebalancing is heavily biased toward the majority class, leading to poor performance on the positive class. Moreover, while using an ensemble of unbalanced SVMs helps improve the performances, the improvement remains limited and it does not significantly mitigate the bias unless resampling is applied beforehand (Ensemble Learning + Resampling). Among all strategies, Resampling and Cost-Sensitive Learning proved to be the most effective in addressing class imbalance, even when combined. In particular, the best result in terms of positive F1 is obtained with Resampling for the hard abandonment task ($F1=0.434$), and with Resampling + Cost-Sensitive Learning for the soft abandonment task ($F1=0.528$). These findings highlight the substantial impact of class imbalance on model performance, reinforcing the importance of carefully selecting strategies to mitigate its effects. The significant performance gap between the model without any resampling technique and the best-performing approaches emphasizes how ignoring class imbalance leads to biased predictions, particularly disadvantaging the minority class. While ensemble methods alone provide only marginal improvements, integrating Resampling and Cost-Sensitive Learning proves to be the most effective way to improve the performance of the model in this task.
9 Discussion
9.1 Detecting abandoning users
@@
Churn prediction has important social and practical implications [79]. Here we defined and tackled the novel task of predicting churn (i.e., user abandonment) following a massive deplatforming moderation intervention on Reddit [18]. In consideration of the novelty of the task —tackled here for the first time— our results are promising. For example, we achieved micro F1 $=0.914$ when detecting users who immediately halted activity after the moderation intervention (hard abandonment), and micro F1 $=0.814$ when detecting users who initially maintained some activity but who eventually left Reddit (soft abandonment). Albeit preliminary, these results demonstrate that the task of preemptively estimating the effects of future moderation interventions is indeed feasible. Adding to the encouraging findings are the results of our leave-one-out cross-validation (LOOCV) analysis, which underscore the absence of substantial performance degradation when applying our model to unseen data and shows that the model learned generalizable behavioral patterns. At the same time however, our study highlighted some of the challenges of this new task. Among them are the difficulties at forecasting complex and heterogeneous user reactions [19], a dire class imbalance [50], a partial and unreliable view of platform moderation actions [4], and the growing difficulties at accessing platform data [80], which culminate in the lack of extensive labeled datasets. The inherent challenges of predicting the effects of moderation interventions —exemplified by our promising yet preliminary results— imply that much work remains to be done before moderators and platform administrators will be able to leverage powerful and dependable tools for accurately estimating the effects of their actions. To this end, this study paves the way for future endeavors in this emerging area of content moderation, which proposes a new pathway for empowering moderators to optimize interventions based on desired outcomes, such as reducing toxicity while minimizing user attrition and churn. This is currently an as-yet unexplored frontier of moderation, which will open up once it achieves high performance in tasks such as the one proposed and tackled in this work.
9.2 Characterizing moderated users
@@
Although the performances of the models developed in this study are promising, there is much room for improvement. Model performance strongly depends on the informativeness of the provided features. In this initial study, we computed and tested an extensive and diverse set of 142 features, many of which were borrowed from recent literature on related tasks [81, 31, 63, 60]. Our set of features included information about the activity of the users before the moderation intervention, their participation in multiple communities, their relationships with other users, their use of toxic or otherwise aggressive speech, and their writing style. While some of the extracted features turned out to be powerful predictors in our models, a large share of the features that we evaluated was not particularly informative, which hindered model performance. This suggests that our task, proposed here for the first time, is substantially different from others previously addressed in related literature and for which those features provided robust performance. This finding underscores the importance of tailoring feature selection to the specific characteristics of the task, rather than relying on already-proposed features. Future efforts towards the prediction of moderation intervention effects should thus consider implementing and experimenting with a broader and even more diversified set of features.
Understanding the characteristics that influence user abandonment following a moderation intervention also provides valuable theoretical insights into online community dynamics. Our feature importance analysis highlights key behavioral patterns that distinguish users who leave the platform from those who remain, shedding light on possible psychological and sociological mechanisms underlying their decisions. As discussed in Section 8.4, the trend of comments in non-banned subreddits is among the most informative features. Users who were already decreasing their activity outside the banned communities before the intervention were more likely to abandon the platform entirely, suggesting that their engagement was closely tied to the restricted communities. This aligns with prior research suggesting that individuals with strong ties to specific groups may struggle to integrate into new ones after sudden disruptions [82, 83]. Additionally, the ratio between the number of non-banned and banned subreddits a user participated in proved highly predictive of abandonment. Users who were more deeply embedded in banned communities, with relatively few interactions elsewhere, were more likely to leave. This supports the idea that having alternative engagement opportunities can act as a buffer against platform departure. Indicators of toxicity also emerged as strong predictors of abandonment. Users with higher toxicity levels were more likely to leave after a ban, potentially reflecting a pattern where users who frequently engage in disruptive behaviors feel alienated or unwilling to conform to platform norms following intervention. This observation aligns with theories in digital sociology that suggest enforcement actions can reinforce disengagement among users who strongly identify with rule-violating behaviors [84]. These findings offer preliminary insights into how moderation interventions shape user behavior, with implications for both theoretical understanding and practical applications.
Based on these observations, promising directions for future work involve the development of socio-psychological features. In fact, predicting the effects of moderation interventions essentially revolves around predicting user behavior. Moreover, it is known that socio-psychological characteristics influence online behaviors, including toxic and aggressive ones [19, 85]. A recent demonstration of the latter can be found in those works who measured a significant improvement in hate speech detection tasks, when incorporating socio-psychological features [86]. By the same token, those features could provide valuable information also in the related task of predicting behavioral changes following a moderation intervention, and particularly for those interventions targeting hateful or toxic users.
9.3 Scalability and Implementation
For real-world deployment, scalability and computational efficiency are key considerations. Our approach leverages traditional machine learning models, which are computationally efficient and significantly less resource-intensive than deep learning-based alternatives. Additionally, many of our features can be precomputed and updated incrementally, reducing the overhead of real-time inference. Notably, our feature importance analysis indicates that only a subset of features is sufficient for accurate predictions, suggesting that models could be further optimized by focusing on the most informative features. Moreover, real-time processing is not a strict requirement, as large-scale moderation interventions (e.g., subreddit bans) are relatively infrequent and do not require continuous, high-speed predictions. Nonetheless, future implementations could explore parallel computing, model compression techniques, or cloud-based deployment to further enhance efficiency.
9.4 Cross-platform Applicability
The application of our predictive approach to estimating the effects of moderation intervention across different online platforms presents both opportunities and challenges. In principle, our approach is general and its success in a new environment primarily depends on data availability and the ability to leverage platform-informative features. However, each platform has unique conventions, user behaviors, and moderation practices, which affect the generalizability of predictive models. Nonetheless, the existence of platforms featuring similar characteristics provide initial avenues for adaptation. For instance, Reddit shared structural and behavioral similarities with Voat, a now-defunct platform that attracted deplatformed communities. YouTube’s dynamics resemble those of BitChute, where content creators rely heavily on subscriber engagement. Likewise, Twitter and Weibo operate as fast-paced, influencer-driven networks, while TikTok and Instagram emphasize visual content and algorithm-driven exposure. These similarities could suggest that models trained on one platform could be transferred to analogous platforms with relatively minor modifications. At the same time however, major differences between platforms must be carefully considered. Platforms might differ not only in affordances and user interaction patterns, but also in their moderation policies [4]. While Reddit primarily enforces moderation at the community level (e.g., subreddit bans)  [21, 18], Twitter/X and Facebook focus more on individual accounts  [87, 88], and YouTube’s interventions often revolve around demonetization or content takedowns  [89, 90]. These differences imply that a system trained to estimate the effects of a certain type of intervention cannot be directly used to predict the effects of a different intervention, even within the same platform. Each intervention type —whether account suspensions, content removals, or demonetization— affects users in distinct ways, requiring models specifically tailored to the nature of the intervention. Ultimately, the heterogeneous landscape of online platforms and their evolving moderation practices pose challenges to cross-platform applicability. Adapting predictive moderation tools requires both scientific and practical efforts to ensure that models remain accurate and relevant in different digital ecosystems. Future research should explore the trade-off between platform-independent and platform-specific features so as to best capture user engagement and retention dynamics, while also considering how shifting moderation strategies affect the accuracy of predictive models.
9.5 Predicting effects: classification, quantification, regression
@@
9.6 Ethical considerations
The use of predictive models in content moderation may raise ethical concerns regarding privacy, bias, and transparency. While our study relies solely on publicly available and anonymized Reddit data, real-world applications must ensure compliance with data protection regulations and implement strict privacy safeguards to avoid over-surveillance of users. Biased predictions are another concern. To mitigate bias, we applied feature selection and rebalancing techniques to prevent skewed predictions. Nonetheless, predictive models can still reflect underlying biases in moderation histories. Future research should audit platform moderation practices for fairness, so as to minimize possible unintended discrimination, and explore fairness-aware machine learning approaches for making predictions. Additionally, the inclusion of potentially sensitive user attributes as features should be carefully weighed against their actual predictive value, ensuring that their use is minimized or avoided whenever possible to uphold fairness and privacy. Transparency is also crucial for maintaining trust in predictive moderation. While our study employs SHAP and PI to analyze feature importance, real-world implementations should prioritize explainable AI methods to ensure even better interpretability. Finally, predictive models should not replace human decision-making but serve as tools to support moderators, ensuring that intervention strategies remain fair and accountable.
9.7 Limitations
@@
This study focuses on a set of Reddit users who experienced a specific form of moderation intervention (i.e., multiple community bans) at a specific time, which may limit the generalizability of our findings. Each platform operates within a unique ecosystem with its own user demographics, interaction norms, and moderation policies. Moreover, user behavior and platform moderation dynamics may change and evolve through time. Therefore, in spite of our efforts at estimating the generalizability of our results, findings derived from a single platform and for a single intervention may not be fully applicable to other platforms or interventions. Furthermore, we focused on a subset of all moderated users: those who were particularly active in the banned communities. As such, our models might exhibit decreased performance if applied to users with markedly different characteristics. Additionally, the specific definitions of hard and soft abandonment that we proposed here might represent a further limitation, given that alternative definitions may yield different interpretations of user behavior and corresponding ground truths for the machine learning tasks. Another limitation of our work stems from the set of features with which we experimented. Despite implementing a relatively large number of features that convey diverse information, many important dimensions of user behavior remain unexplored for this task. Encoding those dimensions in effective machine learning features would boost the performance of the trained models, allowing better results. By the same token, in this first study on predicting the effect of a moderation intervention, we focused on detecting abandoning users. However, effects of a moderation intervention can be defined and investigated in multiple other ways that do not necessarily involve user abandonment, but rather the degree of toxic speech on the platform after the intervention, the extent of polarization, the reliance on factual news, and more. The predictability of these and other moderation effects remains to be assessed.
10 Conclusions and Future Work